OpenStudy (anonymous):

line lm is the midsegment of trapezoid ABCD. AB= x + 8, LM = 4x + 3, and DC = 187. What is the value of x?

OpenStudy (anonymous):



OpenStudy (anonymous):

by the midpoint theorem the eq. should be 4x+3= 1/2 (x+8+187) 2(4x+3) = x+195 8x +6 =x+195 or 8x -x =195 -6 7x = 189 x=189/7 x=27